OpenStudy (anonymous):

how do i solve -3=-5+x?? do i get rid of 5 or x first?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

-3+5=x ===> 2=x ==> x=2, you never get rid of x

OpenStudy (anonymous):

-5+x = -3 x = -3+5 x = 2

OpenStudy (anonymous):

or if u want to get rid then -5+x = -3 -5+x+5= -3 +5 x = 2

OpenStudy (anonymous):

in order to isolate x, you need to add 5 to both sides.

OpenStudy (anonymous):

you may add or you may transfer it to the other side of the equation eith opposite sign